Grade 12Electrostatics

If 1 plate of a parallel plate capacitor has charge +Q & another plate has charge -Q,then the attractive force between the 2 plates of the capacitor is

Shreyoshi
8 Years agoGrade 12

1 Answer

Eshan
8 Years ago
Let the area of each plate beA.

Electric field produced by the positive plate=\dfrac{Q}{2A\epsilon_0}
Henc the attractive force acting on the negatively charged plate due to this electric field is
Q(\dfrac{Q}{2A\epsilon_0})=\dfrac{Q^2}{2A\epsilon_0}